# Measurements of low energy nuclear recoil quenching factors for Na and I recoils in the NaI(Tl) scintillator

### v2

@@ I Introduction
Currently, Weekly Interacting Massive Particles (WIMPs), deemed as one of the most promising dark matter candidates [1, 2], have been extensively explored by numerous experiments, yet with no conclusive findings [3, 4]. This has prompted a shift in focus towards the low-mass dark matter realm, driving increased interest in low energy events [5, 6, 7, 8]. This growing interest is relevant not only to dark matter studies but also to neutrino experiments, especially those involving coherent elastic neutrino-nucleus scattering (CENS) [9]. In these experiments, neutrinos interact with target materials, depositing less than a few keV nuclear recoil energies (keVnr), emphasizing the importance of understanding low energy nuclear recoils.
@@
The NaI(Tl) crystal is widely used in applications such as WIMP dark matter searches [13, 14, 15, 16, 17] and neutrino experiments [12]. Positive signals reported by the DAMA/LIBRA experiment [18, 19, 13] have attracted attention, but interpretations of these signals remain uncertain [20, 11, 21]. Although the COSINE-100 data has ruled out specific interpretations [22], model-independent data analysis is inconclusive [23]. The ANAIS-112 experiment presents inconsistent results with DAMA/LIBRA at about 3 level [14], but a potential difference in QFs is suggested by DAMA/LIBRA [13].
DAMA/LIBRA reported QF measurements using neutrons from a Cf source [24]. The measured responses were compared with the simulated neutron energy spectrum to obtain QF values of 0.3 for sodium and 0.9 for iodine, assuming independence from the energy of the recoiling nuclei. Recently, various QF measurements from refined methods of NaI(Tl) crystals with mono-energetic neutron beams have been reported [25, 26, 27, 28, 29, 30]. In these measurements, the detection of an elastically scattered neutron at a fixed angle relative to the incoming neutron beam direction provides energy-dependent QF values significantly lower than DAMA/LIBRA’s QF results, approximately 0.1 at 10 keVnr for sodium. Although different methods of nuclear recoil calibration may contribute to this discrepancy, the crystal dependence of the QF values is not definitively excluded. Since DAMA/LIBRA reported their annual modulation results with a 0.75 keVee energy threshold [31] corresponding to 2.5 keVnr (assuming 0.3 sodium QF of DAMA/LIBRA [24]), understanding low energy nuclear recoil QF is essential.
@@ II Time of flight (TOF) measurements for incident neutron energy calibration
Mono-energetic neutrons are produced through the deuteron-deuteron nuclear fusion reaction using a DD109 neutron generator (Adelphi Technology, Inc.) [34]. The generator tube was successively shielded by borated polyethylene (40 cm thickness) and high-density polyethylene (40 cm thickness), which complies with safety regulations. Neutrons were emitted through a 3.5-cm-diameter hole in the shield at a 90 angle relative to the deuteron beam. The entire experimental setup was aligned in the direction of the neutron beam.
The incident neutron energy was determined using a He detector positioned in front of the collimated hole when the DD109 generator was initially installed, resulting in a measured neutron beam energy of 2.43 0.03 MeV at an 85 keV deuteron energy. However, for previous QF measurements, a 60 keV deuteron energy was utilized to address higher trigger rates caused by higher neutron and x-ray fluxes at higher deuteron energy. Given the reported near independence of neutron energy from the deuteron energy at the 90 angle [35, 36], the assumption was made that the neutron beam energy remains at 2.43 0.03 MeV even when the deuteron energy is reduced to 60 keV. Nevertheless, due to the volume of the deuteron plate and the NaI(Tl) crystal, there could be a certain angle distribution slightly different from 90. For a systematic understanding of the incident neutron energy, we employed the neutron energy calibration method based on time-of-flight (TOF) measurement.
Two liquid scintillator (LS) detectors were installed using EJ-301 from Eljen Technology, previously used for neutron tagging in QF measurements [28]. Adjacent to the collimated hole of the neutron generator, a one-inch diameter LS detector, named LS1, was situated. Additionally, a second LS detector (LS2) with a three-inch diameter was placed at a 53.1 0.3 relative to the neutron beam direction and 320.0 2.0 cm distance, as depicted in Fig. 1.
@@
where is determined as follow,
Here, A is the atomic mass number of the recoiled nuclei, and is the angle between the neutron beam direction and the scattered neutron direction.
The TOF with a deuteron energy of 100 keV was first measured. To assess the TOF of the neutron and gamma events, we differentiate between the induced particles based on their pulse shapes, taking advantage of the longer decay time characteristic of the neutron events [37]. The pulse shape discrimination (PSD) parameter is defined as a ratio of tail charge to total charge [28, 37]. As illustrated in Fig. 2, coincident neutron events and events are effectively identified based on the PSD parameter and different TOFs. The measured TOF values are found to be 240.8 0.3 ns and 6.8 0.1 ns for neutron and events, respectively. With a distance between the two LS detectors of 320.0 2.0 cm, the TOF for gamma events should be 10.7 0.1 ns. Therefore, the measured TOF is corrected by adding 3.9 0.1 ns. The corrected neutron TOF is then determined to be 244.7 0.3 ns. Using Eq. 2 with Hydrogen interaction in the LS detector, the incident neutron energy is evaluated as 2.48 0.07 MeV, where the uncertainty accounts for the uncertainties in the distance and angle.
The incident neutron energy was measured based on the TOFs of two LS detectors for two more deuteron energies, specifically at 60 keV and 80 keV. The resulting measured neutron energies are 2.32 0.06 MeV and 2.41 0.06 MeV for 60 keV and 80 keV, respectively. All results are presented in Fig. 3. Notably, the incident neutron energies exhibit a clear dependence on the deuteron energies.
@@ III.1 Experimental setup
To maximize light collection of a NaI(Tl) crystal, a small-sized crystal with dimensions of 1.8 cm 1.8 cm 1.4 cm is directly attached to two 3-inch R12669 PMTs, as illustrated in Fig. 4. The crystal is grown from the WIMPScint-III powder, produced by Alpha Spectra Inc., and originates from the same ingot as the COSINE-100 crystal 6 and 7 [38]. Instead of having a quartz window to encapsulate the crystal with copper or aluminum case, only a 1 mm thin optical pad connects the crystal to the PMTs optically. This technique has already been proved to enhance light collection [39, 12] efficiency. The crystal-PMTs assembly is housed in an aluminum casing with a thickness of 5 mm.
@@
Nine EJ-301 LS detectors were deployed surrounding the NaI(Tl) detector for neutron tagging. These LS detectors come in three different sizes: four with a 1-inch diameter, three with a 2-inch diameter, and two with a 3-inch diameter, arranged from the beam direction to outer angles. Specifically, to measure low energy nuclear recoils below 5 keVnr for sodium, one of the 1-inch diameter detectors is placed in the co-linear beam direction at 10.4. The arrangement of the DD109 generator, the NaI(Tl) crystals, and the nine neutron tagging detectors is illustrated in Fig. 4.
@@ III.2.1 NaI(Tl) Crystal Calibration and Light Yield
The electron-equivalent energy scale of the NaI(Tl) crystal is linearly calibrated using a single line of 59.54 keV, achieved through exposure to a Am radioactive source. To estimate the light yield, we first identify a cluster of photoelectrons [41], and then assess the light output using the ratio between the charge of the 59.54 keV events and that of a single photoelectron [38]. In this study, a high light yield of 26.0 0.7 NPE/keVee was obtained, facilitated by the novel crystal encapsulation technique [39]. This represents around a 70% increase compared to the average values of 15 NPE/keVee obtained from COSINE-100 detectors with the previous design [38], enabling exploration below 0.5 keVee energy.
We routinely conducted Am calibration measurements every day to monitor the stability of light yield and correct gain variations. Throughout the approximately one-month measurement period, the detector module consistently maintained a stable and high light yield.
@@ III.2.5 QF extraction
Events having coincidences with the neutron detectors, passing nuclear recoil selections, and clearing noise rejection cuts are applied to measure the QFs of the NaI(Tl) crystal. The selected data for the fixed angle of the neutron tagging detector are compared with simulated energy spectra for various QF values. The GEANT4-based simulation includes mono-energetic neutron generation from the deuteron plate of the DD109 generator, the flight of neutrons to the NaI(Tl) crystal, scattering, and tagging of the scattered neutron with the neutron tagging detectors, with the geometry of the neutron measurement setup shown in Fig. 4. Efficiencies from the trigger and event selection are evaluated using waveform simulation data, as illustrated in Fig.7, and are applied to the simulated energy spectra. The evaluated efficiency is about 0.3 at 0.2 keVee (5 NPE) and reaches unity above 15 NPE, corresponding to about 0.6 keVee. The nuclear recoil energy from the collinear angle of 10.4 is evaluated as 3.8 0.6 keVnr. Minimization of between the measured energy spectrum and the simulated spectra from various QF values is performed. The best fit value is obtained from the minimum , and the uncertainty is derived from the =1 range (inset) as 11.2 0.7 % as shown in Fig. 7.
We extract the QF values for both sodium and iodine from nine neutron tagging detectors. Due to relatively low energy nuclear recoils and low QF values, iodine QFs are evaluated for scattering angles above 33.0. Table 1 summarizes the measured QFs of sodium and iodine for nine neutron tagging detectors at different angles. Measured QFs strongly correlate with the nuclear recoil energy, with smaller values for lower recoil energies. The iodine QF is considerably smaller than for sodium recoils, similar to previous reports [25, 28].
@@
| angle () | (keVnr) | QF (%) | (keVnr) | QF (%) |
@@
IV Revisit the previous QF measurement
The decision to revisit the previous QF measurement, reported in Ref.[28], was prompted by the identification of an incident neutron energy calibration issue, as detailed in Sec.II. During this revisitation, we employ a methodology similar to that used for the new measurement, incorporating advanced event selection and efficiency evaluation based on waveform simulation.
The incident neutron energy from the previous measurement has been re-evaluated and adjusted to 2.34 0.06 MeV, differing from the initially reported value of 2.43 0.03 MeV, as depicted in Fig. 3. This revision leads to a lower neutron energy and consequently, a smaller nuclear recoil energy in the NaI(Tl) crystals. For instance, at the lowest scattering angle of the previous measurement (13.2), the originally reported nuclear recoil energy of 5.8 0.4 keVnr has been corrected to 5.5 1.1 keVnr.
@@ IV Revisit the previous QF measurement
In the previous measurement, we evaluated the trigger efficiency using an independent setup with NaI(Tl) and LaBr crystals irradiated by multiple -rays from a Na source [28]. The radioactive decay of Na emits three -rays, including one at 1274 keV and two at 511 keV. The 511 keV from the LaBr crystal was tagged, and low energy events in the NaI(Tl) crystals were identified as coincident Compton scattering events from the other -rays. The measured trigger efficiency was estimated by comparing the total events to the triggered events. However, our assumption that all events in the NaI(Tl) crystals coincident with the 511 keV in the LaBr crystal were pure scintillation events overlooked the possibility of direct hits on the PMT glasses by rays, which could generate fast-decaying Cherenkov light. Analysis of the NaI(Tl) and LaBr setup data revealed a sharp increase in the event rate below 1 keVee, even though 511 keV coincident was required in the LaBr crystal. Because of the asymmetric behavior of the PMT-induced noise events, the trigger probabilities of these events were significantly lower, resulting in a tendency toward lower trigger efficiency, as illustrated in Fig. 8.
@@
The trigger efficiency was calculated based on the waveform simulation and applied to the QF estimation. For validation of the waveform simulation data, the trigger efficiency was reassessed using the same Na calibration data of the NaI(Tl) and LaBr setup. PMT-induced noise events were eliminated from the sample, and the increased random event rate from energetic source irradiation was accounted for by excluding the contribution from random coincidence events. As depicted in Fig.8, the trigger efficiency derived from the waveform simulation aligns well with the reevaluated efficiency. However, the trigger efficiency applied in the previous analysis was markedly lower than that of the revisited analysis, as well as the waveform simulation data.
@@
Figure 8 illustrates the energy spectra of nuclear recoil events tagged by the neutron detector at 16.4, corresponding to 8.7 1.3 keVnr, selected using the previous criteria (blue-solid line) and the revisited criteria (red-solid line). Additionally, the energy spectra corrected by the evaluated efficiencies for both the previous analysis (blue-circle points) and the revisited analysis (red-square points) are presented. Due to the biased trigger efficiency in the previous analysis, the QF value was lower, measuring at 9.6 1.6%. However, after revisiting the analysis method, the corrected QF is 13.0 2.2%.
@@
Table 2 summarizes revisited QF results in comparison to the previous measurements. The application of advanced event selection has led to the recovery of a few small scattering angles, corresponding to lower recoil energies. Slight increases in QFs with decreased incident neutron energy from new calibration are observed for high scattering angles. Revisiting the trigger efficiency at small scattering angles significantly increases the measured QF values. These updated results align with the findings from the new measurement of the high light yield crystal, as well as with measurements by other groups [26, 29, 25] as shown in Fig. 9.
@@
| angle () | Prev. | New | Prev. | New | |

### v3

# Measurements of low-energy nuclear recoil quenching factors for Na and I recoils in the NaI(Tl) scintillator

@@ I Introduction
Currently, weakly interacting massive particles (WIMPs), deemed as one of the most promising dark matter candidates [1, 2], have been extensively explored by numerous experiments, yet with no conclusive findings [3, 4]. This has prompted a shift in focus towards the low-mass dark matter realm, driving increased interest in low energy events [5, 6, 7, 8]. This growing interest is relevant not only to dark matter studies but also to neutrino experiments, especially those involving coherent elastic neutrino-nucleus scattering (CENS) [9]. In these experiments, neutrinos interact with target materials, depositing less than a few keV nuclear recoil energies (keVnr), emphasizing the importance of understanding low energy nuclear recoils.
@@
The NaI(Tl) crystal is widely used in applications such as WIMP dark matter searches [13, 14, 15, 16, 17] and neutrino experiments [12]. Positive signals reported by the DAMA/LIBRA experiment [18, 19, 13] have attracted attention, but interpretations of these signals remain uncertain [20, 11, 21]. Although the COSINE-100 data have ruled out specific interpretations [22], model-independent data analysis is inconclusive [23]. The ANAIS-112 experiment presents inconsistent results with DAMA/LIBRA at about 3 level [14], but a potential difference in QFs is suggested by DAMA/LIBRA [13].
DAMA/LIBRA reported QF measurements using neutrons from a 252Cf source [24]. The measured responses were compared with the simulated neutron energy spectrum to obtain QF values of 0.3 for sodium and 0.09 for iodine, assuming independence from the energy of the recoiling nuclei. Recently, various QF measurements from refined methods of NaI(Tl) crystals with mono-energetic neutron beams have been reported [25, 26, 27, 28, 29, 30]. In these measurements, the detection of an elastically scattered neutron at a fixed angle relative to the incoming neutron beam direction provides energy-dependent QF values significantly lower than DAMA/LIBRA’s QF results, approximately 0.1 at 10 keVnr for sodium. Although different methods of nuclear recoil calibration may contribute to this discrepancy, the crystal dependence of the QF values is not definitively excluded. Since DAMA/LIBRA reported their annual modulation results with a 0.75 keVee energy threshold [31] corresponding to 2.5 keVnr (assuming 0.3 sodium QF of DAMA/LIBRA [24]), understanding low energy nuclear recoil QF is essential.
@@ II Time of flight (TOF) measurements for incident neutron energy calibration
Mono-energetic neutrons are produced through the deuteron-deuteron nuclear fusion reaction using a DD109 neutron generator (Adelphi Technology, Inc.) [34]. The generator tube was successively shielded by borated polyethylene (40 cm thickness) and high-density polyethylene (40 cm thickness), which complies with safety regulations. Neutrons were emitted through a 3.5-cm-diameter hole in the shield at a 90∘ angle relative to the deuteron beam. The entire experimental setup was aligned in the direction of the neutron beam.
The incident neutron energy was determined using a 3He detector [35, 36] positioned in front of the collimated hole when the DD109 generator was initially installed, resulting in a measured neutron beam energy of 2.43 0.03 MeV at an 85 keV deuteron energy. However, for previous QF measurements, a 60 keV deuteron energy was utilized to address higher trigger rates caused by higher neutron and x-ray fluxes at higher deuteron energy. Given the reported near independence of neutron energy from the deuteron energy at the 90∘ angle [37, 38], the assumption was made that the neutron beam energy remains at 2.43 0.03 MeV even when the deuteron energy is reduced to 60 keV. Nevertheless, due to the volume of the titanium target inside the generator and the NaI(Tl) crystal [34], there could be a certain angle distribution slightly different from 90∘. For a systematic understanding of the incident neutron energy, we employed the neutron energy calibration method based on time-of-flight (TOF) measurement.
Two liquid scintillator (LS) detectors were installed using EJ-301 from Eljen Technology, previously used for neutron tagging in QF measurements [28]. Adjacent to the collimated hole of the neutron generator, a one-inch diameter LS detector, named LS1, was situated. Additionally, a second LS detector (LS2) with a three-inch diameter was placed at a 53.1 0.3∘ relative to the neutron beam direction and 320.0 2.0 cm distance, as depicted in Fig. 1.
@@
where is determined as follows:
Here, is the atomic mass number of the recoiled nuclei, and is the angle between the neutron beam direction and the scattered neutron direction.
The TOF with a deuteron energy of 100 keV was first measured. To assess the TOF of the neutron and events, we differentiate between the induced particles based on their pulse shapes, taking advantage of the longer decay time characteristic of the neutron events [39]. The pulse shape discrimination (PSD) parameter is defined as a ratio of tail charge to total charge [28, 39]. As illustrated in Fig. 2, coincident neutron events and events are effectively identified based on the PSD parameter and different TOFs. The measured TOF values are found to be 240.8 0.3 ns and 6.8 0.1 ns for neutron and events, respectively. With a distance between the two LS detectors of 320.0 2.0 cm, the TOF for events should be 10.7 0.1 ns. Therefore, the measured TOF is corrected by adding 3.9 0.1 ns. The corrected neutron TOF is then determined to be 244.7 0.3 ns. Using Eq. 2 with hydrogen interaction in the LS detector, the incident neutron energy is evaluated as 2.48 0.07 MeV, where the uncertainty accounts for the uncertainties in the distance and angle.
The incident neutron energy was measured based on the TOFs of two LS detectors for two more deuteron energies, specifically at 60 keV and 80 keV. The resulting measured neutron energies are 2.32 0.06 MeV and 2.41 0.06 MeV for 60 keV and 80 keV, respectively. All results are presented in Fig. 3. Notably, the incident neutron energies exhibit a dependence on the deuteron energies.
@@ III.1 Experimental setup
To maximize light collection of a NaI(Tl) crystal, a small-sized crystal with dimensions of 1.8 cm 1.8 cm 1.4 cm is directly attached to two 3-in. R12669 PMTs (Hamamatsu Photonics), as illustrated in Fig. 4. The crystal is grown from the WIMPScint-III powder, produced by Alpha Spectra Inc., and originates from the same ingot as the COSINE-100 crystal 6 and 7 [40]. Instead of having a quartz window to encapsulate the crystal with copper or aluminum case, only a 1 mm thin optical pad EJ-560 (Eljen Technology) connects the crystal to the PMTs optically. This technique has already been proved to enhance light collection [41, 12] efficiency. The crystal-PMTs assembly is housed in an aluminum casing with a thickness of 5 mm.
@@
Nine EJ-301 LS detectors were deployed surrounding the NaI(Tl) detector for neutron tagging. These LS detectors come in three different sizes: four with a 1-in. diameter, three with a 2-in. diameter, and two with a 3-in. diameter, arranged from the beam direction to outer angles. Specifically, to measure low energy nuclear recoils below 5 keVnr for sodium, one of the 1-in. diameter detectors is placed in the colinear beam direction at 10.4∘. The arrangement of the DD109 generator, the NaI(Tl) crystals, and the nine neutron tagging detectors is illustrated in Fig. 4.
@@ III.2.1 NaI(Tl) crystal calibration and light yield
The electron-equivalent energy scale of the NaI(Tl) crystal is linearly calibrated using a single line of 59.54 keV, achieved through exposure to a 241Am radioactive source. To estimate the light yield, we first identify a cluster of photoelectrons [43], and then assess the light output using the ratio between the charge of the 59.54 keV events and that of a single photoelectron [40]. In this study, a high light yield of 26.0 0.7 NPE/keVee was obtained, facilitated by the novel crystal encapsulation technique [41]. This represents around a 70% increase compared to the average values of 15 NPE/keVee obtained from COSINE-100 detectors with the previous design [40], enabling exploration below 0.5 keVee energy.
We routinely conducted 241Am calibration measurements every day to monitor the stability of light yield and correct gain variations. Throughout the approximately one-month measurement period, the detector module consistently maintained a stable and high light yield.
@@ III.2.5 QF extraction
Events having coincidences with the neutron detectors, passing nuclear recoil selections, and clearing noise rejection cuts are applied to measure the QFs of the NaI(Tl) crystal. The selected data for the fixed angle of the neutron tagging detector are compared with simulated energy spectra for various QF values. The GEANT4-based simulation includes mono-energetic neutron generation from the deuteron plate of the DD109 generator, the flight of neutrons to the NaI(Tl) crystal, scattering, and tagging of the scattered neutron with the neutron tagging detectors, with the geometry of the neutron measurement setup shown in Fig. 4. Efficiencies from the trigger and event selection are evaluated using waveform simulation data, as illustrated in Fig.7, and are applied to the simulated energy spectra. The evaluated efficiency is about 0.3 at 0.2 keVee (5 NPE) and reaches unity above 15 NPE, corresponding to about 0.6 keVee. The nuclear recoil energy from the collinear angle of 10.4∘ is evaluated as 3.8 0.6 keVnr. Minimization of between the measured energy spectrum and the simulated spectra from various QF values is performed. The best fit value is obtained from the minimum , and the uncertainty is derived from the = 1 range (inset) as 11.2 0.7 % as shown in Fig. 7. The relative event rates between the data and the simulation at each scattering angle consistently agree, indicating no potential miscalculation of QF due to low energy event selection [45].
We extract the QF values for both sodium and iodine from nine neutron tagging detectors. Due to relatively low energy nuclear recoils and low QF values, iodine QFs are evaluated for scattering angles above 33.0∘. Table 1 summarizes the measured QFs of sodium and iodine for nine neutron tagging detectors at different angles. Measured QFs strongly correlate with the nuclear recoil energy, with smaller values for lower recoil energies. The iodine QF is considerably smaller than for sodium recoils, similar to previous reports [25, 28].
@@
| angle (∘) | (keVnr) | QF (%) | (keVnr) | QF (%) |
@@
IV Re-examine the previous QF measurement
The decision to reexamine the previous QF measurement, reported in Ref. [28], was prompted by the identification of an incident neutron energy calibration issue, as detailed in Sec. II. During this reexamination, we employ a methodology similar to that used for the new measurement, incorporating advanced event selection and efficiency evaluation based on waveform simulation.
The incident neutron energy from the previous measurement has been re-evaluated and adjusted to 2.34 0.06 MeV, differing from the initially reported value of 2.43 0.03 MeV, as depicted in Fig. 3. This revision leads to a lower neutron energy and consequently, a smaller nuclear recoil energy in the NaI(Tl) crystals. For instance, at the lowest scattering angle of the previous measurement (13.2∘), the originally reported nuclear recoil energy of 5.8 0.4 keVnr has been corrected to 5.5 1.1 keVnr.
@@ IV Re-examine the previous QF measurement
In the previous measurement, we evaluated the trigger efficiency using an independent setup with NaI(Tl) and LaBr3 crystals irradiated by multiple rays from a 22Na source [28]. The radioactive decay of 22Na emits three rays, including one at 1274 keV and two at 511 keV. The 511 keV from the LaBr3 crystal was tagged, and low energy events in the NaI(Tl) crystals were identified as coincident Compton scattering events from the other rays. The measured trigger efficiency was estimated by comparing the total events to the triggered events. However, our assumption that all events in the NaI(Tl) crystals coincident with the 511 keV in the LaBr3 crystal were pure scintillation events overlooked the possibility of direct hits on the PMT glasses by rays, which could generate fast-decaying Cherenkov light. Analysis of the NaI(Tl) and LaBr3 setup data revealed a sharp increase in the event rate below 1 keVee, even though 511 keV coincident was required in the LaBr3 crystal. Because of the asymmetric behavior of the PMT-induced noise events, the trigger probabilities of these events were significantly lower, resulting in a tendency toward lower trigger efficiency, as illustrated in Fig. 8.
@@
The trigger efficiency was calculated based on the waveform simulation and applied to the QF estimation. For validation of the waveform simulation data, the trigger efficiency was reassessed using the same 22Na calibration data of the NaI(Tl) and LaBr3 setup. PMT-induced noise events were eliminated from the sample, and the increased random event rate from energetic source irradiation was accounted for by excluding the contribution from random coincidence events. As depicted in Fig.8, the trigger efficiency derived from the waveform simulation aligns well with the reevaluated efficiency. However, the trigger efficiency applied in the previous analysis was markedly lower than that of the re-examined analysis, as well as the waveform simulation data.
@@
Figure 8 illustrates the energy spectra of nuclear recoil events tagged by the neutron detector at 16.4∘, corresponding to 8.7 1.3 keVnr, selected using the previous criteria (blue-solid line) and the re-examined criteria (red-solid line). Additionally, the energy spectra corrected by the evaluated efficiencies for both the previous analysis (blue-circle points) and the re-examined analysis (red-square points) are presented. Due to the biased trigger efficiency in the previous analysis, the QF value was lower, measuring at 9.6 1.6%. However, after reexamining the analysis method, the corrected QF is 13.0 2.2%.
@@
Table 2 summarizes re-examined QF results in comparison to the previous measurements. The application of advanced event selection has led to the recovery of a few small scattering angles, corresponding to lower recoil energies. Slight increases in QFs with decreased incident neutron energy from new calibration are observed for high scattering angles. Re-examining the trigger efficiency at small scattering angles significantly increases the measured QF values. These updated results align with the findings from the new measurement of the high light yield crystal, as well as with measurements by other groups [26, 29, 25] as shown in Fig. 9.
@@
| angle (∘) | Prev. | New | Prev. | New | |